# Parameter control for eccentric, precessing binary black hole simulations with SpEC

### v1

@@ Abstract
Numerical relativity simulations of merging black holes provide the most accurate description of the binary dynamics and the emitted gravitational wave signal. However, practical considerations such as imperfect initial data and initial parameters mean that achieving target parameters, such as the orbital eccentricity or the black hole spin directions, at the beginning of the usable part of the simulation is challenging. In this paper, we devise a method to produce simulations with specific target parameters, namely the Keplerian orbital parameters - eccentricity, semi-major axis, mean anomaly - and the black hole spin vectors using SpEC. The method is an extension of the current process for achieving vanishing eccentricity and it is based on a parameter control loop that iteratively numerically evolves the system, fits the orbit with analytical post-Newtonian equations, and calculates updated input parameters. Through SpEC numerical simulations, we demonstrate $\lesssim 10^{-3}$ and ${\cal{O}}(\rm degree)$ convergence for the orbital eccentricity and the spin directions respectively in $\leq 7$ iterations. These tests extend to binaries with mass ratios $q\leq 3$, eccentricities $e\leq 0.65$, and spin magnitudes $|\chi|\leq 0.75$. Our method for controlling the orbital and spin parameters of numerical simulations can be used to produce targeted simulations in sparsely covered regions of the parameter space or study the dynamics of relativistic binaries.
@@ I Introduction
NR solves an initial value problem where initial conditions are evolved forward in time. For binary BH (BBH) systems, one can freely specify the ratio of the BH masses, the initial BH spin angular momenta (magnitude and direction), and the initial coordinate positions and velocities of the BHs. For puncture data Brandt and Bruegmann (1997) these are indeed directly the free parameters. For quasi-equilibrium conformal thin sandwich data Pfeiffer and York (2003); Cook and Pfeiffer (2004); Caudill et al. (2006) —which is the type of initial-data considered here, as it provides initial lapse and shift for evolutions and allows for nearly extremal black holes Lovelace et al. (2008)— the positions and velocities of the BHs are instead encoded by the binary’s orbital velocity, the BH separation, and the BH relative radial velocity. Evolving the initial data forward in time yields the full spacetime dynamics, BH properties, and emitted gravitational wave signal as a function of time.
However, in practice the initial data for a BBH simulation do not correspond to a snapshot of a binary that has been inspiraling since the infinite past. This is because the initial data do not include the correct gravitational radiation that was emitted in the past during the infinitely-long inspiral, nor do they include the appropriate tidal distortions of the BHs. As a result, when an NR simulation begins to evolve, initial transients occur before the BBH relaxes into a quasiequilibrium state. The gravitational radiation emitted during this process is known as “junk radiation”. Junk radiation, if it is resolved by the simulation, is a physical solution of the Einstein equations, but is not the solution of interest. There have been multiple attempts to reduce the amount of junk radiation in NR simulations Alvi (2000); Yunes and Tichy (2006); Johnson-McDaniel et al. (2009); Kelly et al. (2010); Reifenberger and Tichy (2012); Tichy (2017); Lovelace (2009); Varma et al. (2018), but currently the standard practice is to simply discard the first few orbits until the junk radiation has decayed away Boyle et al. (2019); Pretto et al. (2024).
@@
Among orbital parameters, the eccentricity is particularly astrophysically relevant. Since gravitational radiation reaction efficiently circularizes BBH orbits Peters (1964), most simulations have targeted vanishing eccentricities. Motivated by this, the SpEC workflow includes an initial “eccentricity removal” stage that tunes the input parameters, i.e., the angular velocity, separation, and relative radial velocity, to reduce eccentricity below a predetermined threshold Buonanno et al. (2011). An iterative scheme makes initial guesses for the input parameters, performs an NR simulation for a few orbits, measures the eccentricity, and then chooses updated input parameters for the next iteration. However, recent hints of nonzero eccentricity in select observed signals Romero-Shaw et al. (2021, 2022); Ramos-Buades et al. (2023b); Gupte et al. (2024) and an emphasis of eccentricity as a tool to study the BBH formation history Zevin et al. (2021) have reinvigorated interest in eccentric BBHs Romero-Shaw et al. (2023); Ramos-Buades et al. (2022); Nagar et al. (2024).
@@
In this paper we extend the eccentricity removal procedure of Ref. Buonanno et al. (2011) to non-zero eccentricity; our method also specifies the spin directions at some chosen time different than $t=0$. After introducing the problem setup, in Sec. II we describe the method. Given the target parameters of the simulation we would like to evolve, i.e., target Keplerian parameters and spins, we iteratively evolve the system, extract the binary parameters at the usable part of the data, and adjust the input parameters. Each evolution proceeds for 3-5 orbits, at the end of which we fit the numerical data with a post-Newtonian model and extract the Keplerian parameters. We then use standard root-finding techniques to compute new input parameters that when evolved will result in a simulation that is closer to the target parameters.
@@ II Initial conditions for Numerical Relativity
Consider a binary of two BHs with masses $m_{A}$, with $A\in\{1,2\}$, total mass $M\equiv m_{1}+m_{2}$ , mass ratio $q\equiv m_{1}/m_{2}\geq 1$, and symmetric mass ratio $\eta=q/(1+q)^{2}$. The initial dimensionless spins of the BHs are denoted $\vec{\chi}_{A}$. To simulate such a binary, SpEC requires initial data parameters $\vec{\theta}_{\rm ID}=\{\Omega_{0},v_{r,0},D_{0}\}$ where $\Omega_{0}$ is the orbital angular frequency, $v_{r,0}$ is the coordinate relative radial velocity, and $D_{0}$ is the coordinate BH separation at simulation time $t=0$. Other SpEC papers adopt the convention $\dot{a}_{0}=v_{r,0}/D_{0}$ instead of $v_{r,0}$, e.g. Buonanno et al. (2011); Ossokine et al. (2015). We use $v_{r,0}$ here and reserve the letter $a$ for the semi-major axis of elliptical orbits.
As the evolution proceeds, the coordinate trajectories are described by the BH positions $\vec{x}_{A}(t)$. If the orbit were Newtonian, the trajectory could be equivalently described by a Keplerian parametrization in terms of orbital elements $\vec{\theta}_{\rm orb}=\{a,e,\ell\}$, where $a$ is the semi-major axis, $e$ is the eccentricity, and $\ell$ is the mean anomaly at some reference time. The actual trajectories are however not Newtonian, but rather post-Newtonian (PN) with an increasing number of relevant corrections as the BHs approach merger Boyle et al. (2007). As we consider a small portion of the whole trajectory far from merger, namely the first few orbital periods, and we still approximately parameterize the orbit by $\vec{\theta}_{\rm orb}$. We describe this parametrization and correspondence between $\vec{\theta}_{\rm orb}$ and $\vec{\theta}_{\rm ID}$ in Sec. II.4.
@@
The goal is to devise a method of choosing appropriate input values of the initial data parameters $\vec{\theta}_{{\rm ID},T}$ and the initial spins $\vec{\chi}_{A,T}(t=0)$ so that we obtain simulations with target orbital parameters $\vec{\theta}_{\rm orb}{}_{,T}$ and spins $\vec{\chi}_{A,T}(t=t_{T})$ at some time $t_{T}$. Throughout the paper, we use the subscript $T$ to denote the “target” parameters, including both the orbital parameters $\vec{\theta}_{\rm orb}{}_{,T}$ and the corresponding initial data parameters $\vec{\theta}_{{\rm ID},T}$ that when evolved will give an orbit with $\vec{\theta}_{\rm orb}{}_{,T}$ (and similarly for the spins). Figure 1 presents the problem setup in schematic form. Input parameters $\vec{\theta}_{\rm ID}$ and initial spins $\vec{\chi}_{A}(t=0)$ are set at the initial time $t=0$ and the system is numerically evolved (blue). The usable portion of the numerical data starts after the end of the junk radiation phase (grey band), denoted at $t_{T}$ (black dotted vertical line). At that time, the spins $\tilde{\chi}_{A}(t=t_{T})$ are read off the numerical data, while $\vec{\theta}_{\rm orb}$ is obtained as described in Sec. II.4.
When specifying the target spins $\vec{\chi}_{A,T}(t=t_{T})$ at the target time $t_{T}$, a suitable reference frame needs to be selected. By convention, SpEC simulations begin at $t=0$ with a frame whose $z$-axis is aligned with the orbital angular momentum and the BHs are on the $x$-axis, with the larger BH on the positive $x$-axis. During the simulation, the BHs move in these inertial coordinates, and the simulation outputs the BH inertial frame components of the spins. At time $t_{T}$ –at which we want to control the BH spins, the BHs will in general not be on the inertial coordinate x-axis, and will in general not move in the xy-plane of the inertial coordinates. In order to specify target spins independent of the precise inertial frame position of the black holes, we introduce the BH frame, which is coorbiting in the sense that its x-axis is always pointing from BH 1 to BH 2, and its z-axis is orthogonal to the instantaneous orbital plane. (At $t=0$, the BH frame coincides with the inertial frame). For clarity, we will denote spins in the BH frame with a vector, e.g., the target spin is $\vec{\chi}_{A,T}(t=t_{T})$, and spins in the inertial frame (which is the frame the NR data express them in) with a tilde.
@@
    Due to junk radiation, we have to set $t_{T}\gtrsim 100\,$M. During that time, spin-precession changes the BH spin directions in the inertial simulation frame. This change happens on the (slow) precession timescale. In the co-orbiting BH frame where we define the target spins, the spins evolve on the much faster orbital timescale. To minimize impact of this fast time-scale, we select $t_{T}$ to correspond to an integer number of orbits, after which the BH frame has roughly returned to its initial position. Here, we choose $t_{T}$ to be one orbit. This choice is explained further in Sec. II.5.
@@
$\vec{\chi}_{A}^{\,(i)}(t=0)$ Initial spin vector for BH A for the $i^{\rm th}$ trial simulation
$\tilde{\chi}_{A}^{\,(i)}(t=t_{T})$ Spin vector for BH A at $t=t_{T}$ in the inertial frame for the $i^{\rm th}$ trial simulation, read off the NR data
@@
$\vec{\theta}_{\rm ID}^{\,(i)}$ Initial data parameters for the $i^{\rm th}$ trial simulation
$\vec{\theta}_{\rm orb}^{\,(i)}$ Orbital parameters for the $i^{\rm th}$ trial simulation
@@ II.1 The Inverse Problem and Root-Finding
The map between the input parameters $\vec{\theta}_{\rm ID}$ at $t=0$ and the orbital parameters $\vec{\theta}_{\rm orb}$ at a later time $t_{T}$ is written as an operator, ${\cal{S}}$ which includes:
@@
Collectively, these steps are denoted as, ${\cal{S}}(\vec{\theta}_{\rm ID})$ in Fig. 1. The inverse mapping is denoted by ${\cal{G}}$ in Fig. 1.
For some target orbital parameters, $\vec{\theta}_{\rm orb}{}_{,T}$, we seek $\vec{\theta}_{\rm ID,\,T}$ such that:
@@
Suppose we begin with an initial data guess $\vec{\theta}_{\rm ID}^{\,(i)}$ that yields orbital parameters $\vec{\theta}_{\rm orb}^{\,(i)}$. The superscript here denotes the $i^{\rm th}$ trial evolution as part of an iterative scheme. If we update $\vec{\theta}_{\rm ID}^{\,(i)}$ to $\vec{\theta}_{\rm ID}^{\,(i+1)}=\vec{\theta}_{\rm ID}^{\,(i)}+\delta\vec{\theta}_{\rm ID}$, a Taylor expansion yields
@@
If we knew the operator $\cal G$ exactly, then the update would converge in one iteration using the procedure above. Unfortunately, the exact $\cal G$ is unknown. Therefore, the primary purpose of this paper is to derive suitable approximations to $\cal G$ such that the iterative procedure converges in a reasonable number of iterations. In Sec. II.2 we describe the iterative process, while in Sec. II.3 we use post-Newtonian Keplerian equations to approximate the map between $\vec{\theta}_{\rm orb}$ and $\vec{\theta}_{\rm ID}$.
@@ II.2 Parameter control loop
We adopt the eccentricity threshold imposed by the initial implementation of the parameter control loop in SpEC aiming at vanishing eccentricity Buonanno et al. (2011): $|e_{T}-e^{(i)}|\leq 7\times 10^{-4}$. Even though we only set a tolerance threshold on the eccentricity, in practice we find that all orbital and spin parameters converge to their target value reasonably well. Adding further tolerance thresholds on spins or further orbital parameters is straightforward. Having established the iterative parameter update procedure, we turn to the post-Newtonian equations required for the fits and updates of Boxes 4 and 6.

### v2

@@ Abstract
Numerical relativity simulations of merging black holes provide the most accurate description of the binary dynamics and the emitted gravitational wave signal. However, practical considerations such as imperfect initial data and initial parameters mean that achieving target parameters, such as the orbital eccentricity or the black hole spin directions, at the beginning of the usable part of the simulation is challenging. In this paper, we devise a method to produce simulations with specific target parameters, namely the Keplerian orbital parameters—eccentricity, semimajor axis, mean anomaly—and the black hole spin vectors using SpEC. The method is an extension of the current process for achieving vanishing eccentricity and it is based on a parameter control loop that iteratively numerically evolves the system, fits the orbit with analytical post-Newtonian equations, and calculates updated input parameters. Through SpEC numerical simulations, we demonstrate $\lesssim 10^{-3}$ and ${\cal{O}}(\rm degree)$ convergence for the orbital eccentricity and the spin directions respectively in $\leq 7$ iterations. These tests extend to binaries with mass ratios $q\leq 3$, eccentricities $e\leq 0.65$, and spin magnitudes $|\chi|\leq 0.75$. Our method for controlling the orbital and spin parameters of numerical simulations can be used to produce targeted simulations in sparsely covered regions of the parameter space or study the dynamics of relativistic binaries.
@@ I Introduction
NR solves an initial value problem where initial conditions are evolved forward in time. For binary BH (BBH) systems, one can freely specify the ratio of the BH masses, the initial BH spin angular momenta (magnitude and direction), and the initial coordinate positions and velocities of the BHs. For puncture data Brandt and Bruegmann (1997) these are indeed directly the free parameters. For quasiequilibrium conformal thin sandwich data Pfeiffer and York (2003); Cook and Pfeiffer (2004); Caudill et al. (2006) —which is the type of initial-data considered here, as it provides initial lapse and shift for evolutions and allows for nearly extremal black holes Lovelace et al. (2008)— the positions and velocities of the BHs are instead encoded by the binary’s orbital velocity, the BH separation, and the BH relative radial velocity. Evolving the initial data forward in time yields the full spacetime dynamics, BH properties, and emitted gravitational wave signal as a function of time.
However, in practice the initial data for a BBH simulation do not correspond to a snapshot of a binary that has been inspiraling since the infinite past. This is because the initial data do not include the correct gravitational radiation that was emitted in the past during the infinitely long inspiral, nor do they include the appropriate tidal distortions of the BHs. As a result, when an NR simulation begins to evolve, initial transients occur before the BBH relaxes into a quasiequilibrium state. The gravitational radiation emitted during this process is known as ”junk radiation” Junk radiation, if it is resolved by the simulation, is a physical solution of the Einstein equations, but is not the solution of interest. There have been multiple attempts to reduce the amount of junk radiation in NR simulations Alvi (2000); Yunes and Tichy (2006); Johnson-McDaniel et al. (2009); Kelly et al. (2010); Reifenberger and Tichy (2012); Tichy (2017); Lovelace (2009); Varma et al. (2018); Etienne (2024), but currently the standard practice is to simply discard the first few orbits until the junk radiation has decayed away Boyle et al. (2019); Pretto et al. (2024).
@@
Among orbital parameters, the eccentricity is particularly astrophysically relevant. Since gravitational radiation reaction efficiently circularizes BBH orbits Peters (1964), most simulations have targeted vanishing eccentricities. Motivated by this, the SpEC workflow includes an initial “eccentricity removal” stage that tunes the input parameters, i.e., the angular velocity, separation, and relative radial velocity, to reduce eccentricity below a predetermined threshold Buonanno et al. (2011); Habib et al. (2024). An iterative scheme makes initial guesses for the input parameters, performs an NR simulation for a few orbits, measures the eccentricity, and then chooses updated input parameters for the next iteration. However, recent hints of nonzero eccentricity in select observed signals Romero-Shaw et al. (2021, 2022); Ramos-Buades et al. (2023b); Gupte et al. (2024) and an emphasis of eccentricity as a tool to study the BBH formation history Zevin et al. (2021) have reinvigorated interest in eccentric BBHs Romero-Shaw et al. (2023); Ramos-Buades et al. (2022); Nagar et al. (2024).
@@
In this paper we extend the eccentricity removal procedure of Ref. Buonanno et al. (2011) to nonzero eccentricity; our method also specifies the spin directions at some chosen time different than $t=0$. After introducing the problem setup, in Sec. II we describe the method. Given the target parameters of the simulation we would like to evolve, i.e., target Keplerian parameters and spins, we iteratively evolve the system, extract the binary parameters at the usable part of the data, and adjust the input parameters. Each evolution proceeds for 3-5 orbits, at the end of which we fit the numerical data with a post-Newtonian model and extract the Keplerian parameters. We then use standard root-finding techniques to compute new input parameters that when evolved will result in a simulation that is closer to the target parameters.
@@ II Initial conditions for Numerical Relativity
Consider a binary of two BHs with masses $m_{A}$, with $A\in\{1,2\}$, total mass $M\equiv m_{1}+m_{2}$ , mass ratio $q\equiv m_{1}/m_{2}\geq 1$, and symmetric mass ratio $\eta=q/(1+q)^{2}$. The initial dimensionless spins of the BHs are denoted $\vec{\chi}_{A}$. To simulate such a binary, SpEC requires initial data parameters $\vec{\theta}_{\rm ID}=\{\Omega_{0},v_{r,0},D_{0}\}$ where $\Omega_{0}$ is the orbital angular frequency, $v_{r,0}$ is the coordinate relative radial velocity, and $D_{0}$ is the coordinate BH separation at simulation time $t=0$. Other SpEC papers adopt the convention $\dot{a}_{0}=v_{r,0}/D_{0}$ instead of $v_{r,0}$, e.g. Buonanno et al. (2011); Ossokine et al. (2015). We use $v_{r,0}$ here and reserve the letter $a$ for the semimajor axis of elliptical orbits.
As the evolution proceeds, the coordinate trajectories are described by the BH positions $\vec{x}_{A}(t)$. If the orbit were Newtonian, the trajectory could be equivalently described by a Keplerian parametrization in terms of orbital elements $\vec{\theta}_{\rm orb}=\{a,e,\ell\}$, where $a$ is the semimajor axis, $e$ is the eccentricity, and $\ell$ is the mean anomaly at some reference time. The actual trajectories are however not Newtonian, but rather post-Newtonian (PN) with an increasing number of relevant corrections as the BHs approach merger Boyle et al. (2007). As we consider a small portion of the whole trajectory far from merger, namely the first few orbital periods, and we still approximately parametrize the orbit by $\vec{\theta}_{\rm orb}$. We describe this parametrization and correspondence between $\vec{\theta}_{\rm orb}$ and $\vec{\theta}_{\rm ID}$ in Sec. II.4.
@@
The goal is to devise a method of choosing appropriate input values of the initial data parameters $\vec{\theta}_{{\rm ID},T}$ and the initial spins $\vec{\chi}_{A,T}(t=0)$ so that we obtain simulations with target orbital parameters $\vec{\theta}_{\rm orb}{}_{,T}$ and spins $\vec{\chi}_{A,T}(t=t_{T})$ at some time $t_{T}$. Throughout the paper, we use the subscript $T$ to denote the “target” parameters, including both the orbital parameters $\vec{\theta}_{\rm orb}{}_{,T}$ and the corresponding initial data parameters $\vec{\theta}_{{\rm ID},T}$ that when evolved will give an orbit with $\vec{\theta}_{\rm orb}{}_{,T}$ (and similarly for the spins). Figure 1 presents the problem setup in schematic form. Input parameters $\vec{\theta}_{\rm ID}$ and initial spins $\vec{\chi}_{A}(t=0)$ are set at the initial time $t=0$ and the system is numerically evolved (blue). The usable portion of the numerical data starts after the end of the junk radiation phase (gray band), denoted at $t_{T}$ (black dotted vertical line). At that time, the spins $\tilde{\chi}_{A}(t=t_{T})$ are read off the numerical data, while $\vec{\theta}_{\rm orb}$ is obtained as described in Sec. II.4.
When specifying the target spins $\vec{\chi}_{A,T}(t=t_{T})$ at the target time $t_{T}$, a suitable reference frame needs to be selected. By convention, SpEC simulations begin at $t=0$ with a frame whose $z$-axis is aligned with the orbital angular momentum and the BHs are on the $x$-axis, with the larger BH on the positive $x$-axis. During the simulation, the BHs move in these inertial coordinates, and the simulation outputs the BH inertial frame components of the spins. At time $t_{T}$—at which we want to set the BH spins—the BHs will in general not be on the inertial coordinate x-axis, and will in general not move in the xy-plane of the inertial coordinates. In order to specify target spins independent of the precise inertial frame position of the black holes, we introduce the BH frame, which is coorbiting in the sense that its x-axis is always pointing from BH 1 to BH 2, and its z-axis is orthogonal to the instantaneous orbital plane. (At $t=0$, the BH frame coincides with the inertial frame). For clarity, we will denote spins in the BH frame with a vector, e.g., the target spin is $\vec{\chi}_{A,T}(t=t_{T})$, and spins in the inertial frame (which is the frame the NR data express them in) with a tilde.
@@
    Due to junk radiation, we have to set $t_{T}\gtrsim 100\,$M. During that time, spin-precession changes the BH spin directions in the inertial simulation frame. This change happens on the (slow) precession timescale. In the co-orbiting BH frame where we define the target spins, the spins evolve on the much faster orbital timescale. To minimize impact of this fast timescale, we select $t_{T}$ to correspond to an integer number of orbits, after which the BH frame has roughly returned to its initial position. Here, we choose $t_{T}$ to be one orbit. This choice is explained further in Sec. II.5.
@@
$\vec{\chi}_{A}^{\,(i)}(t=0)$ Initial spin vector for BH A for the $i$th trial simulation
$\tilde{\chi}_{A}^{\,(i)}(t=t_{T})$ Spin vector for BH A at $t=t_{T}$ in the inertial frame for the $i$th trial simulation, read off the NR data
@@
$\vec{\theta}_{\rm ID}^{\,(i)}$ Initial data parameters for the $i$th trial simulation
$\vec{\theta}_{\rm orb}^{\,(i)}$ Orbital parameters for the $i$th trial simulation
@@ II.1 The inverse problem and root-finding
The map between the input parameters $\vec{\theta}_{\rm ID}$ at $t=0$ and the orbital parameters $\vec{\theta}_{\rm orb}$ at a later time $t_{T}$ is written as an operator, ${\cal{S}}$ which includes
@@
Collectively, these steps are denoted as, ${\cal{S}}(\vec{\theta}_{\rm ID})$ in Fig. 1. The inverse mapping is denoted by ${\cal{G}}$ in Fig. 1. While ${\cal{S}}(\vec{\theta}_{\rm ID})$ can be fully calculated with SpEC, $\mathcal{G}$ is unknown, which we will elaborate upon later.
For some target orbital parameters, $\vec{\theta}_{\rm orb}{}_{,T}$, we seek $\vec{\theta}_{\rm ID,\,T}$ such that
@@
Suppose we begin with an initial data guess $\vec{\theta}_{\rm ID}^{\,(i)}$ that yields orbital parameters $\vec{\theta}_{\rm orb}^{\,(i)}$. The superscript here denotes the $i$th trial evolution as part of an iterative scheme. If we update $\vec{\theta}_{\rm ID}^{\,(i)}$ to $\vec{\theta}_{\rm ID}^{\,(i+1)}=\vec{\theta}_{\rm ID}^{\,(i)}+\delta\vec{\theta}_{\rm ID}$, a Taylor expansion yields
@@
If we knew the operator $\cal G$ exactly, then the update would converge in one iteration using the procedure above. Unfortunately, the exact $\cal G$ is unknown as it corresponds to the inverse operation of a full numerical simulation. Therefore, the primary purpose of this paper is to derive suitable approximations to $\cal G$ such that the iterative procedure converges in a reasonable number of iterations. In Sec. II.2 we describe the iterative process, while in Sec. II.3 we use post-Newtonian Keplerian equations to approximate the map between $\vec{\theta}_{\rm orb}$ and $\vec{\theta}_{\rm ID}$.
@@ II.2 Parameter control loop
For simplicity, we adopt the eccentricity threshold imposed by the initial implementation of the parameter control loop in SpEC aiming at vanishing eccentricity Buonanno et al. (2011): $|e_{T}-e^{(i)}|\leq 7\times 10^{-4}$. This tolerance is set just below the eccentricity value for which spin-induced oscillations dominate the eccentricity-induced oscillations in $\dot{\Omega}$. Therefore, it is a sufficient constraint for a parameter control loop that relies on $\dot{\Omega}$.
Even though we only set a tolerance threshold on the eccentricity, in practice we find that all orbital and spin parameters converge to their target value reasonably well. Adding further tolerance thresholds on spins or further orbital parameters is straightforward. Having established the iterative parameter update procedure, we turn to the post-Newtonian equations required for the fits and updates of boxes 4 and 6.